OpenStudy (anonymous):

The curve cuts the y axis at (0,2) and the x axis at (-1,0),(1,0) and (2,0) What is the equation of the curve?

ganeshie8 (ganeshie8):

basically you're give 3 zeroes/roots of the polynomial and the value at 0

ganeshie8 (ganeshie8):

do u knw how to write a polynomial wid zeroes of -1, 1, 2 ?

OpenStudy (anonymous):

No don't think so haha

OpenStudy (anonymous):

\[(x-1)(x+1)(x-2)=f(x)\]

ganeshie8 (ganeshie8):

if k is a zero of polynomial, then (x-k) is a factor

OpenStudy (anonymous):

Still don't get it

OpenStudy (kc_kennylau):

if you have a factored polynomial with factor (x-a), then a is one of the zeros. Now you're just reversing the process.

ganeshie8 (ganeshie8):

since \(-1, 1, 2\) are zeroes, \((x--1), (x-1), (x-2)\) will be factors. so the polynomial can be represented as :- \(f(x) = k(x+1)(x-1)(x-2)\) you need to find \(k\) and plugin above

OpenStudy (anonymous):

The question as asked has many solutions. The question should specify that the function is a polynomial.